The car is going 30s speed of 20m/s then with the speed of 10m/s it drives 300m. more whats the middle speed of the car.

1
Expert's answer
2021-12-09T10:08:54-0500

The average speed:


"s=\frac{d}{t}=\frac{30·20+300}{30+300/10}=15\text{ m/s}."